OpenStudy (anonymous):

Solve the following systems of linear equations using elimination or substitution 1. 2x+5y=-5 and x+3y=3 2. 4x+4y+-16 and 2x+8y=10 3. -x-5y=-13 and 2x+2y=-6 4, 2x+y=-1 and -4x+6y=6 Please help

OpenStudy (anonymous):

for substitution you want to solve for x or y in one of the equations (doesn't matter which) and then plug that equation into the other equation for example: 2x+5y=-5 and x+3y=3 x=3-3y so 2(3-3y)=5y=-5 and solve for y when you find y, you can plug it into x+3y=3 to find x for elimination you want to eliminate one of the variables. for example: 2x+y=-1 and -4x+6y=6 multiply 2x+y=-1 by 6 12x+6y=-6 and then subtract -4x+6y=6 BUT REMEMBER TO DISTRIBUTE THE MINUS SIGN  so you get 16x=-12 and solve for x you can plug what you get for x into 2x+y=1 to find y

OpenStudy (anonymous):

hopefully you can do the others on your own now :)